OpenStudy (anonymous):

express as a single logarithm 4logbm - logbn

OpenStudy (anonymous):

If you mean,\[4\log _ b m-\log _ b n\]then you'd have,\[4\log _ b m-\log _ b n = \log _ b m^4-\log_b n=\log_b \frac{m^4}{n}\]

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

np